A 45 kg patient requires a 450mg dose of a particular drug. A solution contains 150mg of the drug per teaspoon. How may ml of the solution should be given to the patient to deliver the required dose of the drug?
Answer #1

ANSWER.

* Calculation of required dose of the drug.

         Formula = Desired dose/Dose in hand *Quantity.

Here

        Desired dose(D) = 450 mg.

        Dose in hand(H) = 150 mg.

            Quantity (Q) = 5 ml (1Teaspoon).

As per above formula

       Amount of drug required = D/H * Q

                                               = 450 / 150 * 5

                                               = 3 * 5

                                               = 15 ml.